### v1

# Gyroscopically stabilized quantum spin rotors

Abstract
Recent experiments demonstrate all-electric spinning of levitated nanodiamonds with embedded nitrogen-vacancy spins. Here, we argue that such gyroscopically stabilized spin rotors offer a promising platform for probing and exploiting quantum spin-rotation coupling of particles hosting a single spin degree of freedom. Specifically, we derive the effective Hamiltonian describing how an embedded spin affects the rotation of rapidly revolving quantum rotors due to the Einstein-de Haas and Barnett effects, which we use to devise experimental protocols for observing this coupling in state-of-the-art experiments. This will open the door for future exploitations of quantum spin rotors for superposition experiments with massive objects.
The free rotation of a magnetized body can differ significantly from that of the same object if it were non-magnetic. This, perhaps surprising, discrepancy is due to the fact that magnetic moments carry a spin angular momentum that contributes to the total angular momentum of the magnetized body, as quantified by the material-specific gyromagnetic ratio. Remarkably, it is the total angular momentum, i.e. the sum of mechanical and spin angular momentum, that enters the equations of motion of the body [1]. This gives rise to the so-called spin-rotation coupling [2] between rigid-body rotations and magnetic degrees of freedom, as famously witnessed in the Einstein-de Haas and Barnett effects [3, 4, 5, 6, 7]. The impact of spin-rotation coupling typically decreases with particle size, while it is still relevant for microscale and nanoscale magnetic objects [8, 9, 10, 11, 12, 13].
Nanoscale particles of various shapes and with sizes ranging from hundreds of nanometers to tens of microns can be stably suspended in ultra-high vacuum using optical, electrical, or magnetic traps [14]. Monitoring and controlling their linear and rotational motion with laser fields enables precision sensing of forces and torques as required for fundamental tests of physics [15, 16]. Recent demonstrations of cooling their center-of-mass oscillations [17, 18, 19, 20] and their angular librations [21, 22] in an optical tweezer into the quantum regime constitute the first steps towards quantum applications of levitated particles. The demonstrated degree of control and read-out renders these systems ideal for performing quantum experiments with objects of unrivaled masses [23, 24, 25, 26, 27, 28].
Figure 1: (a) A charged spheroidal nanodiamond with an embedded nitrogen-vacancy (NV) center is levitated in a Paul trap and revolves rapidly with frequency $\omega$ in the symmetry plane of the trap. The major principal axis $\mathbf{n}_{1}$ of the body is chosen to be aligned with the NV quantization axis so that the NV spin states $|S_{1}=\pm\hbar\rangle$ are split by the rotation-induced Barnett field as well as by an applied magnetic field $\mathbf{B}$. The particle rotation induces (b) an effective gyroscopic potential $I\omega^{2}/2\sin^{2}\beta$ stabilizing the out-of-plane rotation angle $\beta$, up to small librations $\xi$, while the external magnetic field together with spin-rotation coupling induce (c) a spin state-dependent potential $V_{m}(\gamma)$ for $\gamma$ rotations around the body-fixed symmetry axis depending on the spin states $|S_{1}=\hbar m\rangle$ with $m\in\{-1,0,1\}$.
Experiments with levitated magnetic objects aim at harnessing their magnetic properties to gain better control over the internal and external degrees of freedom of the nanoparticle [29]. Magnetized particles have been trapped in electric [30, 31], magnetic [32, 33, 34], optical [35, 36], and superconducting traps [37, 38], where their center-of-mass and rotational motion can be monitored and manipulated. Various schemes have been devised to control their mechanical motion via driving the internal magnetization field [39, 40, 41, 42, 43] or individual magnetic defect centers, such as nitrogen-vacancy (NV) centers [44, 45, 29, 46, 47]. These schemes promise to enable sensing of magnetic fields and accelerations [48, 9] as well as the creation of massive superpositions [49, 26, 50, 51], and perhaps even probing the quantumness of gravity [52].
In this article, we show how the spin-rotation coupling between a levitated nanoparticle and a single embedded spin degree of freedom can be observed and exploited for quantum experiments. The proposed scheme is based on the recent experimental demonstration of rapid spinning of electrically levitated nanodiamonds with embedded spin degrees of freedom [53, 31]. This spinning boosts the spin-rotation coupling while it also gyroscopically stabilizes the nanoparticle orientation in the co-rotating frame (see Fig. 1), so that a combination of magneto-static fields and microwave pulses then enables controlling the collective motion of hundreds of millions of atoms via a single spin.
Specifically, we consider a symmetric nanodiamond with a single embedded NV spin levitated in the presence of an external static magnetic field $\mathbf{B}=B\mathbf{e}_{z}$, as schematically depicted in Fig. 1. The nanodiamond is approximated as an ellipsoidal rigid rotor with semi-axis lengths $l_{1}=l_{2}<l_{3}$, and moments of inertia $I_{1}=I_{2}=I$ with $I_{3}<I$. The nanoparticle orientation $\Omega=(\alpha,\beta,\gamma)$, parametrized by Euler angles in the $z$-$y^{\prime}$-$z^{\prime\prime}$ convention, specifies the tensor ${\rm R}(\Omega)$ which rotates from the lab-frame axis ${\bf e}_{j}$, $j=x,y,z$, to the rotor principal axes $\mathbf{n}_{\mu}(\Omega)={\rm R}(\Omega){\bf e}_{j}$, $\mu=1,2,3$. For instance ${\bf n}_{1}(\Omega)={\rm R}(\Omega){\bf e}_{x}$, etc.; an explicit expression for ${\bf n}_{\mu}(\Omega)$ is given in App. A. The inertia tensor depends on the orientation of the particle through $\mathrm{I}(\Omega)=\sum_{\mu}I_{\mu}\mathbf{n}_{\mu}(\Omega)\otimes\mathbf{n}_{\mu}(\Omega)$. A single NV center with spin-one angular momentum vector operator $\mathbf{S}$ is embedded in the nanodiamond, effectively introducing an internal spin triplet and adding the magnetic dipole moment $\gamma_{0}{\bf S}$ to the particle (with gyromagnetic ratio $\gamma_{0}>0$). The spin angular momentum adds to the mechanical angular momentum to yield the total angular momentum ${\bf J}$ of the system. As the spin quantization axis we choose $\mathbf{n}_{1}(\Omega)$, effectively breaking the symmetry of the particle shape; the NV zero-field splitting is $D_{\rm nv}\simeq 2\pi\times 2.87$ GHz.
Gyroscopic stabilization is achieved by trapping a charged particle in an azimuthally symmetric quadrupole trap and spinning it with high rotation frequency $\omega$ around the lab-frame ${\bf e}_{z}$-axis, as recently realized in two experiments [53, 31]. For homogeneously charged particles the electric quadrupole tensor aligns with the inertia tensor, with quadrupole moments $Q_{1}=Q_{2}=Q$ and $Q_{3}>Q$. For vanishing electric dipole moment, the center-of-mass and rotational motion of the particle decouple so that the resulting rotational macromotion potential is determined by the ac voltage $U_{\rm ac}$, the electrode separation $d_{0}$, and the ac frequency $\omega_{\rm ac}$ [54]. The spin-rotor Hamiltonian is
$H=$ $\dfrac{(\mathbf{J}-\mathbf{S})^{2}}{2I}+\left(\dfrac{1}{2I_{3}}-\dfrac{1}{2I}\right)(J_{3}-S_{3})^{2}+\frac{D_{\mathrm{nv}}}{\hbar}S^{2}_{1}$
$-\gamma_{0}B\left(S_{1}\sin\beta\cos\gamma-S_{2}\sin\beta\sin\gamma-S_{3}\cos\beta\right)$
$+\dfrac{U_{\mathrm{ac}}^{2}(Q-Q_{3})^{2}}{16I\omega_{\mathrm{ac}}^{2}d_{0}^{4}}\sin^{2}\beta\cos^{2}\beta.$ (1)
Here, $J_{\mu}={\bf J}\cdot{\bf n}_{\mu}(\Omega)$ denotes the body-frame component of the total angular momentum vector and likewise for ${\bf S}$. The confining quadrupole potential depends only on the polar angle $\beta$, implying that the lab-frame angular momentum component $p_{\alpha}={\bf J}\cdot{\bf e}_{z}$ is conserved. It is thus convenient to transform the spin-rotor state $|\psi\rangle$ to a co-rotating frame of boosted angular momentum via
$|\psi\rangle=\exp\left(i\frac{I\omega}{\hbar}\alpha-i\omega t\frac{p_{\alpha}}{\hbar}\right)|\phi\rangle.$ (2)
This adds the gyroscopic potential $I\omega^{2}/2\sin^{2}\beta$ [see Fig. 1(b) and App. B] to the quadrupole potential in Eq. (Gyroscopically stabilized quantum spin rotors), effectively confining the particle to the rotation plane $\beta\simeq\pi/2$ with harmonic frequency
$\omega_{\xi}=\sqrt{\omega^{2}+\dfrac{U_{\mathrm{ac}}^{2}(Q-Q_{3})^{2}}{8I^{2}\omega_{\mathrm{ac}}^{2}d_{0}^{4}}}.$ (3)
For large rotation rates $\omega_{\xi}\simeq\omega$, librations out of the rotation plane can be first linearized in $\xi=\pi/2-\beta$ and then adiabatically eliminated. Specifically, linearization yields
$H_{\rm rot}=$ $H_{\xi}+\frac{p^{2}_{\gamma}}{2I_{3}}+\frac{D_{\rm nv}}{\hbar}S_{1}^{2}+(\gamma_{0}BS_{3}-\omega p_{\gamma})\xi$
$+\left(\omega-\gamma_{0}B\right)(S_{1}\cos\gamma-S_{2}\sin\gamma),$ (4)
with $p_{\gamma}=J_{3}$, $H_{\xi}=p_{\xi}^{2}/2I+I\omega_{\xi}^{2}\xi^{2}/2$ and we neglected non-dominant contributions to spin-rotation coupling as well as dispersion in the gyroscopic motion (see App. B). The fast librations out of the rotation plane can now be eliminated by neglecting transitions between eigenstates $|n\rangle$ of $H_{\xi}$ in a Born-Oppenheimer approximation (see App. C). This is achieved by expanding
$|\phi\rangle=\sum_{n=0}^{\infty}\exp\left[-\frac{ip_{\xi}}{\hbar}\frac{\omega p_{\gamma}-\gamma_{0}BS_{3}}{I\omega^{2}}\right]e^{-itH_{\xi}/\hbar}|n,\chi_{n}\rangle,$ (5)
so that the adiabatic spin-rotor state $|\chi_{n}\rangle$ evolves according to the Hamiltonian
Figure 2: (a) Steady-state alignment $\langle\cos\gamma\rangle$ as a function of the applied magnetic field. The variance is indicated by the shaded region. (b) Potential energy surfaces $\Omega_{+}$ (solid) and $\Omega_{-}$ (dashed) for $B=-0.5\,\mathrm{mT}$ and $\omega/2\pi=1\,\mathrm{MHz}$. The inset shows the avoided crossing around $\gamma=\pi/2$ where the spin-rotor state $|\chi_{+}\rangle$ can be stably trapped in the $|\sigma_{x}=+1\rangle$ spin state. (c) Spin transition probability for an initial state in spin $|\sigma_{x}=+1\rangle$ (solid) and spin $|\sigma_{x}=-1\rangle$ (dashed) as a function of $\omega_{\gamma}t$ for different magnetic field strengths. The nanodiamond has an ellipsoidal shape with semi-axes $l_{3}=200$ nm and $l_{1}=l_{2}=0.3l_{3}$.
$H_{n}=$ $\dfrac{p_{\gamma}^{2}}{2I_{\rm eff}}+\dfrac{\gamma_{0}B}{I\omega}S_{3}p_{\gamma}+\dfrac{D_{\mathrm{nv}}}{2\hbar}(S_{1}^{2}+S_{2}^{2})$
$+\frac{1}{2}(\omega-\gamma_{0}B)\left[f_{n}(S_{1}-iS_{2})e^{-i\gamma}+{\rm h.c.}\right],$
$+\dfrac{D_{\mathrm{nv}}}{4\hbar}\left[g_{n}(S_{1}+iS_{2})^{2}+{\rm h.c.}\right].$ (6)
This Hamiltonian describes the coupling between the NV spin and $\gamma$-rotations with effective inertia $I_{\rm eff}=II_{3}/(I-I_{3})$. The final two terms are the adiabatic potential (or potential energy surface) for fixed libration state $|n\rangle$, containing the magnetic potential as well as spin-rotation coupling and the Barnett effect on the NV zero-field splitting. The adiabatic terms are weighted by the overlap integrals
$f_{n}=$ $\langle n|\exp\left(-\frac{ip_{\xi}}{\hbar}\frac{\hbar(\omega-\gamma_{0}B)}{I\omega^{2}}\right)|n\rangle$ (7a)
$g_{n}=$ $\langle n|\exp\left(-\frac{ip_{\xi}}{\hbar}\frac{2\hbar\gamma_{0}B}{I\omega^{2}}\right)|n\rangle$ (7b)
for transitions between displaced and non-displaced libration states. Note that $f_{n}\simeq 1$ for $n\ll I^{2}\omega^{4}/\hbar^{2}\gamma_{0}^{2}B^{2}$ and likewise $g_{n}\simeq 1$. Given that $\hbar\gamma_{0}B/I\omega^{2}\ll 1$ for realistic magnetic fields and rotation rates, the Hamiltonian (Gyroscopically stabilized quantum spin rotors) becomes approximately independent of $n$. Thus, all adiabatic spin-rotor states $|\chi_{n}\rangle$ evolve according to the same effective Hamiltonian
$H_{\mathrm{eff}}=$ $\dfrac{p_{\gamma}^{2}}{2I_{\rm eff}}+\dfrac{D_{\mathrm{nv}}}{\hbar}S_{1}^{2}+\dfrac{\gamma_{0}B}{\omega}\frac{S_{3}p_{\gamma}}{I}$
$+(\omega-\gamma_{0}B)\left(S_{1}\cos\gamma-S_{2}\sin\gamma\right).$ (8)
This Hamiltonian is the main result of this work and the starting point for the proposals discussed below. It accounts for spin-rotation coupling in three different ways: First, the planar $\gamma$-rotation occurs with the effective moment of inertia $I_{\rm eff}$ because $\xi$-librations are gyroscopically stabilized and adiabatically eliminated. Second, the rotation rate $\omega$ of the particle induces the effective spin-dependent potential in the second line of the Hamiltonian (Gyroscopically stabilized quantum spin rotors). Third, the $\mathbf{B}$ field induces a coupling of the spin to rotations around the symmetry axis of the particle as described by the last term in the first line. Note that this coupling is typically small and will be neglected in the following. Next, we will study the implications of gyroscopic stabilization for the experimental observation of strong spin-rotation coupling.
(a) Barnett splitting— The Hamiltonian shows how the degeneracy of the $|S_{1}=\pm\hbar\rangle$ states is lifted by the Barnett effect for $B=0$. Specifically, neglecting $\gamma$-rotations by setting $\gamma=0$ and $p_{\gamma}=0$, the $|S_{1}=\pm\hbar\rangle$ levels are split by the gyroscopic rotation frequency $\omega$. Note that the resonance frequency is also Doppler shifted by the particle rotation rate since the microwave drive is applied in the lab frame rather than in the rotating frame of the NV center [55, 56, 57].
(b) Barnett anti-alignment— The described setup can be used to observe that the Barnett field $-\omega/\gamma_{0}$ tends to anti-align prolate particles with their own rotation axis [11]. This follows from considering the hard magnet limit $|\omega-\gamma_{0}B|\ll D_{\rm nv}$ where the spin is rigidly attached to the body frame $S_{1}=\hbar m$ with $m\in\{-1,0,1\}$, and $S_{2}=S_{3}=0$. Then the adiabatic Hamiltonian (Gyroscopically stabilized quantum spin rotors) implies that the particle revolves according to the spin-dependent potential $V_{m}(\gamma)=\hbar m(\omega-\gamma_{0}B)\cos\gamma$ [see Fig. 1(c)],
$I_{\rm eff}\ddot{\gamma}=\hbar m(\omega-\gamma_{0}B)\sin\gamma.$ (9)
For $B=0$ and $m=1$, the rotor tends to orient with $\gamma=0$ so that the rotation axis anti-aligns with the spin axis. A finite magnetic field $B>0$ weakens the impact of the Barnett field and exactly compensates it at $B=\omega/\gamma_{0}$, where the particle evolves freely in $\gamma$, as would be expected for an object carrying no spin degrees of freedom. Experimentally observing the anti-alignment of the particle spin axis with the rotation axis is thus a unique signature of strong spin-rotation coupling in a rotating nanoparticle. It can be observed in state-of-the-art experiments with nanodiamonds spinning in quadrupole ion traps. This is illustrated in Fig. 2(a), which shows the steady-state alignment $\langle\cos\gamma\rangle$ of the NV axis as a function of the rotation rate $\omega$ for three different temperatures.
(c) Stabilization of spin superpositions— For moderate magnetic fields, the superposition of the magnetic spin states $|\sigma_{x}=+1\rangle=(|S_{1}=+\hbar\rangle+|S_{1}=-\hbar\rangle)/\sqrt{2}$ becomes stabilized when the spin quantization axis lies in the rotation plane. Specifically, for $|\omega-\gamma_{0}B|\lesssim D_{\rm nv}$, the non-magnetic state $|S_{1}=0\rangle$ remains far detuned from the magnetic states $|S_{1}=\pm\hbar\rangle$ for all orientations $\gamma$. The former can thus be adiabatically eliminated, yielding the spin-rotor Hamiltonian in the magnetic spin subspace
$H_{\mathrm{mag}}$ $=\dfrac{p_{\gamma}^{2}}{2I_{\mathrm{eff}}}+\frac{\hbar\delta}{2}(\mathbb{1}+\sigma_{x})\sin^{2}\gamma+\hbar g\sigma_{z}\cos\gamma.$ (10)
Here, $\delta=(\omega-\gamma_{0}B)^{2}/D_{\rm nv}$, $g=(\omega-\gamma_{0}B)$, and $\sigma_{z}$ and $\sigma_{x}$ are Pauli matrices in the magnetic subspace of the NV spin. In contrast to the hard magnet limit, this Hamiltonian accounts for spin transitions as described by $\delta\neq 0$. Diagonalizing the spin-dependent part of Eq. (10) yields the potential energy surfaces $\hbar\Omega_{\pm}(\gamma)/2$ for the $\gamma$ dynamics
$\Omega_{\pm}(\gamma)=\delta\sin^{2}\gamma\pm\sqrt{\delta^{2}\sin^{4}\gamma+4g^{2}\cos^{2}\gamma},$ (11)
as depicted in Fig. 2(b). They exhibit avoided crossings at $\gamma=\pi/2$ (and $23\pi/$ 2), in addition to the potential minima at $\gamma=0,\pi$ for $|S_{1}=\pm\hbar\rangle$ already observed in the hard magnet limit. The energies can be linearized near the avoided crossing yielding $\hbar\Omega_{\pm}(\gamma)=\pm\hbar\delta/2\pm\hbar g^{2}(\gamma-\pi/2)^{2}/\delta$, showing that the spin-rotor is harmonically trapped in the spin state $|\sigma_{x}=+1\rangle$. The groundstate of the resulting harmonic potential is stably trapped for $|\delta|\gg|g\sigma_{\gamma}|$, where $\sigma_{\gamma}=(\hbar\delta/8I_{\rm eff}g^{2})^{1/4}$ is the groundstate width in the effective harmonic potential (see App. E). The state becomes unstable if this condition is violated, as illustrated in Fig. 2(c) for three different magnetic field strengths, or when the spin superposition decoheres after the $T_{2}$ time.
(d) Spin-controlled quantum interference— For strong magnetic fields, $|\omega-\gamma_{0}B|\simeq D_{\rm nv}$, the quantum state of the NV spin can be used to generate and read out superpositions of the angle $\gamma$. Specifically, the magnetic state $|S_{1}=+\hbar\rangle$ can be adiabatically eliminated for small librations around $\gamma\simeq 0$, restricting the spin to the two levels $\left|\uparrow\right\rangle=|S_{1}=-\hbar\rangle$ and $\left|\downarrow\right\rangle=|S_{1}=0\rangle$ with detuning $\Delta=D_{\rm nv}-g$. For small amplitudes and detunings $\langle\gamma^{2}\rangle\ll|\Delta|/|g|\ll 1$, the two spins are effectively decoupled and the resulting dispersive spin-rotor Hamiltonian follows as (App. F)
$H_{\rm d}=\frac{p_{\gamma}^{2}}{2I_{\mathrm{eff}}}+\dfrac{\hbar g}{8}\gamma^{2}+\dfrac{\hbar\Delta}{2}\sigma_{z}+\frac{3\hbar g}{8}\left(1+\frac{4g}{3\Delta}\right)\gamma^{2}\sigma_{z}.$ (12)
[…]
$\frac{1}{\lambda_{\tau}}=1+(1-\mathrm{e}^{2i\omega_{\gamma}\tau})\sinh^{2}\left(\frac{\sqrt{\hbar}g\tau}{\sqrt{I_{\rm eff}\Delta}}\right).$ (13)
The spin probability rephases for $\tau=\pi/\omega_{\gamma}$, where $\lambda_{\tau}=1$, due to a perfect overlap of the states of the two branches, reduced only by NV spin decoherence with realistic dephasing time of $T_{2}=10\,\mu\mathrm{s}$ [58, 59, 60]. This can be seen in Fig. 3(b), which demonstrates that increasing the rotation frequency $\omega$ enhances the interference signal by reducing the total interferometer time. Fig. 3(c) shows the total interferometer time and the recurrence probability $P_{\uparrow}(\tau=\pi/\omega_{\gamma})$ as a function of the spin-rotor coupling $g/D_{\rm nv}$. The shaded region indicates where the NV’s $T_{2}$ time exceeds the interferometer time, resulting in a significant enhancement of the interference signal.
The proposed experiments for gyroscopically stabilized spin-rotors are within reach given the recent achievements in the field of levitated particles [45, 17, 19, 18, 31, 53]. Electrically charged nanodiamonds hosting NV centers have been trapped, manipulated, and monitored both via light scattering as well as by preparation and measurements of the embedded spins (state initialization with a green laser and microwave driving to the magnetic states) [44, 30, 61]. Gyroscopic spinning of charged particles with embedded spin degrees of freedom was recently demonstrated [31, 53], achieving MHz rotation rates. Illuminating the particle with circularly polarized light can further boost the rotation rate beyond GHz [62, 63, 64].
In summary, we discussed the gyroscopic stabilization of rapidly spinning nanodiamonds with a single embedded NV spin, leading to an effective spin-rotor Hamiltonian containing only a single rotational degree of freedom. We proposed several schemes to use this Hamiltonian to probe spin-rotation coupling in state-of-the-art setups and to perform mechanical quantum interference experiments. Our work presents the first steps towards applications of gyroscopically stabilized spin rotors.
Acknowledgments— We thank Cosimo C. Rusconi, Sanchar Sharma, Maxime Perdriat, Julen Pedernales, and Klaus Hornberger for stimulating discussions and comments on the manuscript. VW and BAS acknowledge funding by the Carl-Zeiss-Foundation through the project QPhoton. VW acknowledges support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) via the Research Training Group 1995 ‘Quantum Many-Body Methods in Condensed Matter Systems’ Project Number 240766775. VW and SVK are supported by the German Federal Ministry of Education and Research (BMBF) project QECHQS (Grant No. 16KIS1590K). BAS is supported by the DFG–510794108.
Appendix A Rotation tensor and spin-rotor Hamiltonian
The rotation tensor $\mathrm{R}(\Omega)$ rotates the laboratory frame axes $\mathbf{e}_{x},\mathbf{e}_{y},\mathbf{e}_{z}$ into the body-fixed principal axes $\mathbf{n}_{1},\mathbf{n}_{2},\mathbf{n}_{3}$
$\mathbf{n}_{\mu}(\Omega)=\mathrm{R}(\Omega)\mathbf{e}_{j}.$ (14)
Choosing Euler angles $\Omega=(\alpha,\beta,\gamma)^{T}$ according to the z-y’-z” convention, we have
$\mathrm{R}(\Omega)=$ $\mathrm{R}_{z}(\alpha)\mathrm{R}_{y}(\beta)\mathrm{R}_{z}(\gamma)=\begin{pmatrix}\cos\alpha&-\sin\alpha&0\\ \sin\alpha&\cos\alpha&0\\ 0&0&1\end{pmatrix}$
$\begin{pmatrix}\cos\beta&0&\sin\beta\\ 0&1&0\\ -\sin\beta&0&\cos\beta\end{pmatrix}\begin{pmatrix}\cos\gamma&-\sin\gamma&0\\ \sin\gamma&\cos\gamma&0\\ 0&0&1\end{pmatrix}.$ (15)
This implies that the principal axes in terms of the Euler angles read
$\mathbf{n}_{1}$ $=\begin{pmatrix}\cos\alpha\cos\beta\cos\gamma-\sin\alpha\sin\gamma\\ \sin\alpha\cos\beta\cos\gamma+\cos\alpha\sin\gamma\\ -\sin\beta\cos\gamma\end{pmatrix},$ (16a)
$\mathbf{n}_{2}$ $=\begin{pmatrix}-\cos\alpha\cos\beta\sin\gamma-\sin\alpha\cos\gamma\\ -\sin\alpha\cos\beta\sin\gamma+\cos\alpha\cos\gamma\\ \sin\beta\sin\gamma\end{pmatrix},$ (16b)
$\mathbf{n}_{3}$ $=\begin{pmatrix}\cos\alpha\sin\beta\\ \sin\alpha\sin\beta\\ \cos\beta\end{pmatrix}.$ (16c)
The Hamiltonian of a rigid rotor with embedded spin in an external magnetic field reads [12]
$H_{0}=\dfrac{1}{2}(\mathbf{J}-\mathbf{S})\cdot\mathrm{I}^{-1}(\mathbf{J}-\mathbf{S})+\dfrac{D_{\mathrm{nv}}}{\hbar}(\mathbf{n}\cdot\mathbf{S})^{2}+\gamma_{0}\mathbf{B}\cdot\mathbf{S}.$ (17)
The first term is the rotational kinetic energy with $\mathbf{J}-{\bf S}$ the kinetic angular momentum vector operator of the particle. The second term describes the preferred alignment of the NV quantization axis with the body-fixed ${\bf n}(\Omega)$ axis. These two terms conserve the total angular momentum ${\bf J}$ so that spin and rotation are effectively coupled as witnessed in the Einstein-de Haas and Barnett effects [3, 4]. The third term describes both the torque exerted by the magnetic field on the mechanical rotation as well as the orientation-dependent precession of the NV spin in the external field. For instance, the energy of the spin eigenstates $|S_{\bf n}=\pm 1\rangle$ of the spin operator $S_{\bf n}={\bf n}(\Omega)\cdot{\bf S}$ is Zeeman-split by the magnetic field by the orientation-dependent amount $2\hbar\gamma_{0}{\bf B}\cdot{\bf n}(\Omega)$.
The components of the total and the spin angular momentum operator follow the canonical angular momentum commutation relations. Denoting the body-frame components with Greek indices, $J_{\mu}=\mathbf{J}\cdot\mathbf{n}_{\mu}(\Omega)$ and $S_{\mu}=\mathbf{S}\cdot\mathbf{n}_{\mu}(\Omega)$, with $\mu=1,2,3$, these commutation relations are [1]
$[J_{\lambda},J_{\mu}]=-i\hbar\varepsilon_{\lambda\mu\nu}J_{\nu},\quad{\rm and}\quad[S_{\lambda},S_{\mu}]=i\hbar\varepsilon_{\lambda\mu\nu}S_{\nu},$ (18a)
as well as
$[J_{\lambda},S_{\mu}]=0.$ (18b)
The components of the total and spin angular momentum vector operators can also be given in the lab frame, $J_{j}=\mathbf{J}\cdot\mathbf{e}_{j}$ and $S_{j}={\bf S}\cdot{\bf e}_{j}$. Then $J_{j}$ and $S_{k}$ do not commute as the corresponding canonical commutation relations are
$[J_{k},J_{\ell}]=i\hbar\varepsilon_{k\ell m}J_{m}\quad{\rm and}\quad[S_{k},S_{\ell}]=i\hbar\varepsilon_{k\ell m}S_{m},$ (19a)
as well as
$[J_{k},S_{\ell}]=i\hbar\varepsilon_{k\ell m}S_{m}.$ (19b)
Note that $[J_{j},J_{\mu}]=0$. Thus the operators $({\bf J}^{2},J_{3},J_{z},{\bf S}^{2},S_{\bf n})$ form a complete set of commuting observables [65].
The canonical momentum operators $p_{\alpha},p_{\beta},p_{\gamma}$ are related to the angular momentum vector operator through the relations
$p_{\alpha}=$ ${\bf J}\cdot{\bf e}_{z}$ (20a)
$p_{\beta}=$ $\frac{1}{2}{\bf J}\cdot\left(-\sin\alpha{\bf e}_{x}+\cos\alpha{\bf e}_{y}\right)+{\rm h.c.}$ (20b)
$p_{\gamma}=$ ${\bf J}\cdot{\bf n}_{3}(\Omega).$ (20c)
This implies that the body-frame components of the angular momentum operator can be expressed through
$J_{1}$ $=p_{\beta}\sin\gamma-p_{\alpha}\csc\beta\cos\gamma+\frac{\cot\beta}{2}\{p_{\gamma},\cos\gamma\},$ (21a)
$J_{2}$ $=p_{\beta}\cos\gamma+p_{\alpha}\csc\beta\sin\gamma-\frac{\cot\beta}{2}\{p_{\gamma},\sin\gamma\},$ (21b)
$J_{3}$ $=p_{\gamma}.$ (21c)
Appendix B Trapping potential in the rotating frame
The dynamics of the spin rotor is typically also affected by trapping torques used to control the nanoparticle motion. These torques do not directly affect the spin state and must be added to the Hamiltonian (17). As the experimentally most relevant scenario, we consider the electric torques exerted by radio frequency electric quadrupole traps. Denoting the drive frequency by $\omega_{\rm ac}$ and the ac voltage by $U_{\rm ac}$, the electric field near the trap center is of quadrupole character,
${\bf E}({\bf r},t)=\frac{U_{\rm ac}}{d_{0}}\cos(\omega_{\rm ac}t)\sum_{j=x,y,z}A_{j}({\bf e}_{j}\cdot{\bf r}){\bf e}_{j},$ (22)
where the coefficients $A_{j}$, with $\sum_{j}A_{j}=0$, determine the field gradient and $d_{0}$ is the characteristic length scale of the trap. Here, we aligned the lab frame with the main trapping axes of the quadrupole trap [54]. In practical situations the surface charge distribution of the ellipsoid is approximately body-fixed and homogeneous, so that the dipole moment can be neglected [53] and there is no coupling between center-of-mass and rotational degrees of freedom [54]. The trapping torques are then characterized by the body-fixed quadrupole tensor
$\mathrm{Q}(\Omega)=\sum_{\mu=1}^{3}Q_{\mu}\mathbf{n}_{\mu}(\Omega)\otimes\mathbf{n}_{\mu}(\Omega),$ (23)
which is symmetric and traceless $\sum_{\mu}Q_{\mu}=0$. If the drive frequency is sufficiently fast in comparison to the particle dynamics, so that the small-amplitude micromotion can be temporally averaged, the large-amplitude macromotion is induced by the secular quadrupole potential [54]
$V_{\mathrm{qu}}(\Omega)=\dfrac{U_{\mathrm{ac}}^{2}}{36\omega_{\mathrm{ac}}^{2}d_{0}^{4}}$ $\sum_{\lambda=1}^{3}\frac{1}{I_{\lambda}}\left[\sum_{j=x,y,z}A_{j}\sum_{\mu,\nu=1}^{3}\varepsilon_{\mu\lambda\nu}Q_{\mu}\right.$
×$\times[{\bf n}_{\mu}(\Omega)\cdot{\bf e}_{j}][{\bf n}_{\nu}(\Omega)\cdot{\bf e}_{j}]\right]^{2}.$ (24)
The quadrupole potential depends on all three Euler angles through the principal axes ${\bf n}_{\mu}(\Omega)$. The resulting trapping torques tend to align the principal axes of the particle ${\bf n}_{\mu}(\Omega)$ with the axes of the lab frame ${\bf e}_{j}$. Full three-dimensional alignment is possible if the particle has three distinct quadrupole moments $Q_{1}\neq Q_{2}\neq Q_{3}$ and if the trap is asymmetric $A_{1}\neq A_{2}\neq A_{3}$.
In the considered case of a symmetric particle and symmetric trap configuration, i.e. $A_{1}=A_{2}=-A_{3}/2$, the quadrupole potential reduces to a potential only confining the $\beta$-angle, namely
$V_{\mathrm{qu}}(\beta)=\dfrac{U_{\mathrm{ac}}^{2}(Q-Q_{3})^{2}}{16I\omega_{\mathrm{ac}}^{2}d_{0}^{4}}\sin^{2}\beta\cos^{2}\beta,$ (25)
and Eq. (17) reduces to the Hamiltonian (Gyroscopically stabilized quantum spin rotors). We move to a frame co-rotating with the rotor $UHU^{\dagger}+i\hbar\partial_{t}UU^{\dagger}$ with the unitary operator
$U(t)=\exp\left(i\dfrac{\omega t}{\hbar}p_{\alpha}-i\dfrac{I\omega}{\hbar}\alpha\right),$ (26)
which adds the term $-\omega p_{\alpha}$ and replaces $p_{\alpha}$ by $p_{\alpha}+I\omega$ while leaving the remaining Hamiltonian unchanged. Using that in the rotating frame $p_{\alpha}=0$, the Hamiltonian reads
$H_{\mathrm{rot}}=$ $\dfrac{p_{\beta}^{2}}{2I}+\dfrac{(I\omega-p_{\gamma}\cos\beta)^{2}}{2I\sin^{2}\beta}-\dfrac{p_{\beta}}{I}(S_{1}\sin\gamma+S_{2}\cos\gamma)$
+$+\left\{\dfrac{I\omega-p_{\gamma}\cos\beta}{2I\sin^{2}\beta},S_{1}\cos\gamma-S_{2}\sin\gamma\right\}+V_{\rm qu}(\beta)$
$-\gamma_{0}B\left(S_{1}\sin\beta\cos\gamma-S_{2}\sin\beta\sin\gamma-S_{3}\cos\beta\right)$
$+\dfrac{S_{1}^{2}+S_{2}^{2}}{2I}+\dfrac{(p_{\gamma}-S_{3})^{2}}{2I_{3}}+\frac{D_{\mathrm{nv}}}{\hbar}S^{2}_{1}.$ (27)
Here, a number of terms can be neglected: First, the terms $(S_{1}^{2}+S_{2}^{2})/2I$ and $S_{3}^{2}/2I_{3}$ are negligible in comparison to the NV zero-field splitting $D_{\rm nv}S_{1}^{2}/\hbar$ and can thus be dropped. Second, the spin-rotation coupling in $\beta$ as described by the last term in the first line is always suppressed compared to the $\beta$ trapping potential as well as to the gyroscopic potential in $\gamma$ and can thus be neglected. Third, the spin-rotation coupling in $\gamma$ as described by $-p_{\gamma}S_{3}/I_{3}$ as well as by the second line is always smaller than the gyroscopic coupling described by the second term in the first line and can thus also be dropped. The simplified Hamiltonian reads
$H_{\mathrm{rot}}=$ $\dfrac{p_{\beta}^{2}}{2I}+\dfrac{(I\omega-p_{\gamma}\cos\beta)^{2}}{2I\sin^{2}\beta}+\dfrac{p_{\gamma}^{2}}{2I_{3}}+\frac{D_{\mathrm{nv}}}{\hbar}S^{2}_{1}+V_{\rm qu}(\beta)$
$-\gamma_{0}B\left(S_{1}\sin\beta\cos\gamma-S_{2}\sin\beta\sin\gamma-S_{3}\cos\beta\right)$
$+\dfrac{\omega}{\sin^{2}\beta}(S_{1}\cos\gamma-S_{2}\sin\gamma).$ (28)
By linearizing around $\beta=\pi/2$ and only keeping terms up to second order in $\xi=\pi/2-\beta$, i.e. $1/\sin^{2}\beta\approx 1+\xi^{2}$, $\cos\beta/\sin^{2}\beta\approx\xi$, $\cos^{2}\beta\sin^{2}\beta\approx\xi^{2}$, as well as neglecting the terms proportional to $S_{1}\xi^{2}$, $S_{2}\xi^{2}$, and $p_{\gamma}^{2}\xi^{2}$, the Hamiltonian is
$H_{\rm rot}\approx$ $\dfrac{p_{\xi}^{2}}{2I}+\dfrac{I}{2}\omega_{\xi}^{2}\xi^{2}+(\gamma_{0}BS_{3}-\omega p_{\gamma})\xi+\frac{p^{2}_{\gamma}}{2I_{3}}+\frac{D_{\rm nv}}{\hbar}S_{1}^{2}$
$+\left(\omega-\gamma_{0}B\right)(S_{1}\cos\gamma-S_{2}\sin\gamma).$ (29)
It can be seen that the particle’s rotation adds the gyroscopic potential $I\omega^{2}\xi^{2}/2$ to the quadrupole potential and the spin-rotation coupling induces an effective potential for $\gamma$.
Appendix C Adiabatic approximation
Figure 4: Dynamics of $\gamma$ for the spin state $|S_{z}=-\hbar\rangle$ as a function of $\omega_{0}t$, comparing the solution of (a) the gyroscopic Hamiltonian (Gyroscopically stabilized quantum spin rotors) (light blue, solid) with the adiabatic Hamiltonian (Gyroscopically stabilized quantum spin rotors) (dark blue, dashed) and (b) the adiabatic Hamiltonian (Gyroscopically stabilized quantum spin rotors) with the dispersive Hamiltonian (12) (light blue, solid). The nanodiamond has an ellipsoidal shape with semi-axes $l_{3}=200$ nm and $l_{1}=l_{2}=0.4l_{3}$, $B=-102\,\mathrm{mT}$, $\omega/2\pi=0.1\,\mathrm{MHz}$, $\omega_{0}=\omega_{\gamma}$ and $\sigma_{\gamma}=\sqrt{\hbar/2I_{\mathrm{eff}}\omega_{\gamma}}$. We approximated the Fock space as a finite space of dimension $d=40$ and chose for both the initial states of the $\beta$- and $\gamma$-oscillator a coherent state $|\alpha_{\beta,\gamma}=0.1\rangle$.
The fast $\xi$-librations in Eq. (Gyroscopically stabilized quantum spin rotors) can now be adiabatically eliminated by expanding an arbitrary spin-rotor state $|\phi\rangle$ in the oscillator eigenbasis $|n\rangle$ with $H_{\xi}|n\rangle=\hbar\omega(n+1/2)|n\rangle$ for $\omega_{\xi}\simeq\omega$. Using the ansatz (5),
$|\phi\rangle=\sum_{n=0}^{\infty}W\mathrm{e}^{-itH_{\xi}/\hbar}|n,\chi_{n}\rangle,$ (30)
with $|n,\chi_{n}\rangle=|n\rangle\otimes|\chi_{n}\rangle$ and
$W=\exp\left(-\frac{ip_{\xi}}{\hbar}\frac{\omega p_{\gamma}-\gamma_{0}BS_{3}}{I\omega^{2}}\right),$ (31)
the states $|\chi_{n}\rangle$ then evolve according to the coupled equations
$i\hbar\partial_{t}|\chi_{n}\rangle=$ $\left[\frac{p_{\gamma}^{2}}{2I_{3}}-\hbar\omega\left(n+\dfrac{1}{2}\right)+h_{nn}\right]|\chi_{n}\rangle$
$+\sum_{m=0\atop{m\neq n}}^{\infty}e^{i\omega(n-m)t}h_{nm}|\chi_{m}\rangle.$ (32)
Here we used that $[p_{\gamma},W]=0$ and defined the coupling operators
$h_{nm}$ $=\langle n|W^{\dagger}\left[H_{\xi}+(\gamma_{0}BS_{3}-\omega p_{\gamma})\xi+\dfrac{D_{\mathrm{nv}}}{\hbar}S_{1}^{2}\right.$
[…]
Transitions between distinct libration states, as described by the second line of Eq. (C), can be neglected within the rotating wave approximation (Born-Oppenheimer approximation), so that only the diagonal coupling operators matter. A direct calculation shows that
$h_{nn}=$ $\hbar\omega\left(n+\dfrac{1}{2}\right)-\dfrac{p_{\gamma}^{2}}{2I}-\dfrac{(\gamma_{0}BS_{3})^{2}}{2I\omega^{2}}+\dfrac{\gamma_{0}BS_{3}p_{\gamma}}{I\omega}$
$+\langle n|V(\gamma,\mathbf{S})|n\rangle$ (34)
with the spin-dependent potential
$V(\gamma,\mathbf{S})=$ $\dfrac{D_{\mathrm{nv}}}{\hbar}\left[S_{1}\cos\left(\dfrac{\gamma_{0}B}{I\omega^{2}}p_{\xi}\right)+S_{2}\sin\left(\dfrac{\gamma_{0}B}{I\omega^{2}}p_{\xi}\right)\right]^{2}$
$+(\omega-\gamma_{0}B)\left[S_{1}\cos\left(\tilde{\gamma}\right)-S_{2}\sin\left(\tilde{\gamma}\right)\right].$ (35)
Here the angle $\gamma$ is shifted by the coupling with $p_{\xi}$, i.e.
$\tilde{\gamma}=\gamma+\dfrac{p_{\xi}}{I\omega}-\dfrac{\gamma_{0}Bp_{\xi}}{I\omega^{2}}.$ (36)
By expressing the cosine and sine functions in terms of exponential functions and defining the overlap integrals (7), the effective Hamiltonian $H_{n}$ governing the dynamics of $|\chi_{n}\rangle$ can be written in the form (Gyroscopically stabilized quantum spin rotors). In Fig. 4(a) we compare the dynamics of $\gamma$ close to $\gamma=0$ for Eq. (Gyroscopically stabilized quantum spin rotors) and Eq. (Gyroscopically stabilized quantum spin rotors).
Appendix D Barnett anti-alignment
We calculate the steady-state alignment $\langle\cos\gamma\rangle$ of the NV axis by considering a Boltzmann probability distribution of $\gamma$, i.e.
$\mathrm{prob}(\gamma)=\dfrac{1}{Z}\exp\left[-\dfrac{V_{m}(\gamma)}{k_{\mathrm{B}}T}\right].$ (37)
Here $k_{\mathrm{B}}$ denotes the Boltzmann constant and $Z$ the partition function given as
$Z=\int_{0}^{2\pi}\exp\left(-m\kappa\cos\gamma\right)=2\pi I_{0}(m\kappa),$ (38)
where we defined $\kappa=\hbar(\omega-\gamma_{0}B)/k_{\mathrm{B}}T$ and $I_{\ell}(\cdot)$ is the $\ell$-th order modified Bessel function of the first kind. The expectation value of $\cos\gamma$ follows as
$\langle\cos\gamma\rangle$ $=-\dfrac{1}{Z}\dfrac{1}{m}\dfrac{\mathrm{\partial Z}}{\partial\kappa}=-\dfrac{I_{1}(\kappa m)}{I_{0}(\kappa m)}.$ (39)
Analogously, $\langle\cos^{2}\gamma\rangle=(\mathrm{\partial}^{2}Z/\partial\kappa^{2})/Zm^{2}$ such that the variance is obtained as
$\mathrm{var}(\cos\gamma)=\frac{1}{2}+\dfrac{I_{2}(\kappa m)}{2I_{0}(\kappa m)}-\left[\dfrac{I_{1}(\kappa m)}{I_{0}(\kappa m)}\right]^{2},$ (40)
implying that $20\leq{\rm var}(\cos\gamma)\leq 1/$ 2.
Appendix E Stabilization of spin superpositions
Figure 5: (a) Occupation probability of the $|S_{1}=0\rangle$ spin state calculated from the exact dynamics given by Eq. (Gyroscopically stabilized quantum spin rotors) (dark blue, dashed) compared to the approximated solution of Eq. (43) (light blue, solid). (b) Comparison of the $\gamma$-dynamics obtained from the adiabatic Hamiltonian (Gyroscopically stabilized quantum spin rotors) (dark blue, dashed) and the magnetic spin Hamiltonian (10) (light blue, solid). The nanodiamond has an ellipsoidal shape with semi-axes $l_{3}=200$ nm and $l_{1}=l_{2}=0.4l_{3}$, $B=-0.5\,\mathrm{mT}$, $\omega/2\pi=0.1\,\mathrm{MHz}$ and $\omega_{0}=\hbar/I_{\mathrm{eff}}$. The spin is initially in $\left|\sigma_{x}=-1\right>$ located at $\gamma=\pi/2$. We simulated the dynamics in the angular momentum basis with dimension $d_{m}=1729$ and rescaled $\delta$ and $g$ by a factor $10^{-8}$ due to computational constraints.
To eliminate the spin zero state from the dynamics, we expand the adiabatic spin-rotor state in the eigenbasis of $S_{1}$ as
$|\chi\rangle=\sum_{s=-1}^{+1}|\chi_{s}\rangle|S_{1}=s\hbar\rangle,$ (41)
so that the Hamiltonian (Gyroscopically stabilized quantum spin rotors) yields the coupled equations
$i\hbar\partial_{t}\left|\chi_{0}\right>=$ $\left[\dfrac{p_{\gamma}^{2}}{2I_{\rm eff}}-\hbar D_{\mathrm{nv}}\right]\left|\chi_{0}\right>$
$-\dfrac{\hbar g}{\sqrt{2}}\sin\gamma\left(\left|\chi_{+1}\right>+\left|\chi_{-1}\right>\right),$ (42a)
$i\hbar\partial_{t}\left|\chi_{+1}\right>=$ $\left[\dfrac{p_{\gamma}^{2}}{2I_{\rm eff}}+\hbar g\cos\gamma\right]\left|\chi_{+1}\right>-\dfrac{\hbar g}{\sqrt{2}}\sin\gamma\left|\chi_{0}\right>,$ (42b)
$i\hbar\partial_{t}\left|\chi_{-1}\right>=$ $\left[\dfrac{p_{\gamma}^{2}}{2I_{\rm eff}}-\hbar g\cos\gamma\right]\left|\chi_{-1}\right>-\dfrac{\hbar g}{\sqrt{2}}\sin\gamma\left|\chi_{0}\right>.$ (42c)
For moderate $B$ fields, $|\omega-\gamma_{0}B|\lesssim D_{\rm nv}$, the non-magnetic state $|S_{1}=0\rangle$ remains far detuned from the magnetic states $|S_{1}=\pm\hbar\rangle$ for all orientations $\gamma$. This implies that $i\hbar\partial_{t}\left|\chi_{0}\right>\approx 0$ for all times, so that
$\left|\chi_{0}\right>\approx-\dfrac{\delta}{\sqrt{2}g}\sin\gamma\left(\left|\chi_{+1}\right>+\left|\chi_{-1}\right>\right),$ (43)
as seen in Fig. 5(a). Inserting this approximation into the equations of motion for $|\chi_{\pm 1}\rangle$ allows us to identify the effective Hamiltonian in the magnetic spin subspace as (10). We compare in Fig. 5(b) the $\gamma$-dynamics around $\gamma=\pi/2$ of the adiabatic Hamiltonian with the derived magnetic Hamiltonian. Taylor expanding this Hamiltonian in linear order of $\eta=\gamma-\pi/2\simeq 0$ gives
$H_{\mathrm{mag}}\simeq\frac{p_{\gamma}^{2}}{2I_{\rm eff}}+\frac{\hbar\delta}{2}\sigma_{x}-\hbar g\sigma_{z}\eta=\frac{p_{\gamma}^{2}}{2I_{\rm eff}}+\hbar\Omega\mathrm{e}^{-i\phi\sigma_{y}}\sigma_{x}\mathrm{e}^{i\phi\sigma_{y}},$ (44)
with the angle
$\phi=\frac{1}{2}{\rm arctan}\left(\frac{2g\eta}{\delta}\right)\simeq\frac{g}{\delta}\eta$ (45)
and the frequency
$\Omega=\frac{1}{2}\sqrt{\delta^{2}+4g^{2}\eta^{2}}\simeq\frac{\delta}{2}+\frac{g^{2}}{\delta}\eta^{2}.$ (46)
Here, we assume that $|g\eta|\ll|\delta|$. This condition remains fulfilled for all times for the harmonic oscillator ground state $|n_{\eta}=0\rangle|\sigma_{x}=+1\rangle$ with harmonic frequency $\omega_{\eta}=(2\hbar g^{2}/I_{\mathrm{eff}}|\delta|)^{1/2}$ if
$\left|\frac{g}{\delta}\left(\frac{2I_{\rm eff}\omega_{\eta}}{\hbar}\right)^{1/2}\right|=\left(\frac{\hbar g^{2}}{8I_{\rm eff}|\delta|^{3}}\right)^{1/4}\ll 1.$ (47)
Appendix F Spin-controlled rotational interference
[…]
$\left|\chi_{+1}\right>\approx\dfrac{g}{\sqrt{2}(D_{\mathrm{nv}}+g)}\left|\chi_{0}\right>.$ (48)
The effective Hamiltonian for $\left|\chi_{\downarrow}=\chi_{-1}\right>$ and $\left|\chi_{\uparrow}=\chi_{0}\right>$ then follows as
$H_{2}=\dfrac{p^{2}_{\gamma}}{2I_{\mathrm{eff}}}+\dfrac{\hbar}{4}(g-\tilde{\delta})\gamma^{2}+\dfrac{\hbar}{2}\left[\Delta+\dfrac{(g+\tilde{\delta})}{2}\gamma^{2}\right]\sigma_{z}-\dfrac{\hbar g}{\sqrt{2}}\sigma_{x}\gamma,$ (49)
where we shifted the system by $-\hbar\Delta/2$ and defined
$\tilde{\delta}=\dfrac{g^{2}}{D_{\mathrm{nv}}+g}.$ (50)
The spin-oscillator interaction can be diagonalized by applying the unitary transformation $QH_{2}Q^{\dagger}$ with
$Q=\mathrm{exp}\left(i\dfrac{\pi}{2}\dfrac{|\mathbf{m}|\mathbf{e}_{z}+\mathbf{m}}{||\mathbf{m}|\mathbf{e}_{z}+\mathbf{m}|}\cdot\bm{\sigma}\right),$ (51)
where
$\mathbf{m}=\left[\dfrac{\Delta}{2}+\dfrac{(g+\tilde{\delta})}{4}\right]\mathbf{e}_{z}-\dfrac{g}{\sqrt{2}}\gamma\mathbf{e}_{x}$ (52)
denotes the vector of the local direction of the spin. The transformed Hamiltonian reads
$H_{\mathrm{disp}}\simeq$ $\dfrac{p^{2}_{\gamma}}{2I_{\mathrm{eff}}}+\dfrac{\hbar(g-\tilde{\delta})}{4}\gamma^{2}$
$+\dfrac{\hbar\Delta}{2}\sigma_{z}+\hbar\left(\dfrac{g+\tilde{\delta}}{4}+\dfrac{g^{2}}{\Delta}\right)\sigma_{z}\gamma^{2},$ (53)
where we approximated the spin dependent potential satisfying the conditions of the dispersive limit, i.e.
[…]
and neglected non-adiabatic corrections related to spin-flip transitions. Considering $|\Delta|\ll|g|$ we can approximate $\tilde{\delta}\approx g/2$ for which Eq. (12) follows. Fig. 4(b) shows the comparison of the $\gamma$-dynamics obtained from this Hamiltonian with the one of Eq. (Gyroscopically stabilized quantum spin rotors) close to $\gamma=0$. From this Hamiltonian the evolution operators for the oscillator state $|\chi_{\uparrow(\downarrow)}\rangle$ are obtained as
$U_{\uparrow}$ $=\exp\left[-\dfrac{it}{\hbar}\left(\dfrac{p^{2}_{\gamma}}{2I_{\mathrm{eff}}}+\dfrac{I_{\mathrm{eff}}}{2}\omega_{\gamma}^{2}\gamma^{2}\right)\right],$ (55a)
$U_{\downarrow}$ $=\exp\left[-\dfrac{it}{\hbar}\left(\dfrac{p^{2}_{\gamma}}{2I_{\mathrm{eff}}}+\dfrac{I_{\mathrm{eff}}}{2}\omega_{\gamma}^{2}\gamma^{2}-\dfrac{\hbar g}{4}\left(1+\dfrac{4g}{\Delta}\right)\gamma^{2}\right)\right].$ (55b)
[…]
$P_{\uparrow\downarrow}=\dfrac{1}{2}\pm\dfrac{1}{4}\mathrm{Tr}\left[U_{\downarrow}^{\dagger}U_{\uparrow}^{\dagger}U_{\downarrow}U_{\uparrow}\rho_{\mathrm{th}}+U_{\uparrow}^{\dagger}U_{\downarrow}^{\dagger}U_{\uparrow}U_{\downarrow}\rho_{\mathrm{th}}\right].$ (56)
In the limit $T\ll\hbar\omega_{\gamma}/k_{\rm B}$, meaning that we only consider the ground state $|n_{0}\rangle$, the trace over the evolved thermal state can be calculated as
$I_{\gamma}$ $=\mathrm{Re}\left[\langle n_{0}|U_{\downarrow}^{\dagger}U_{\uparrow}^{\dagger}U_{\downarrow}U_{\uparrow}|n_{0}\rangle\right]$
$=\mathrm{Re}\left[\left(1+\dfrac{(\Delta+4g)^{2}(1-\mathrm{e}^{2i\tau\omega_{\gamma}})}{8g(\Delta+2g)}\sinh^{2}\zeta\right)^{-\frac{1}{2}}\right].$ (57)
For the second equality the Baker-Campbell-Hausdorff formula of SU(1,1) [66] was used to rewrite the evolution operators (55) [12]. For finite temperatures the expression for $I_{\gamma}$ becomes more complicated and can be found in the Supplement of Ref. [12].